# Random divergence-free drifts and the Onsager-Richardson threshold

### 12 Mar 2026 v1

@@ Abstract.
We prove the absence of anomalous dissipation for passive scalars driven by random divergence-free autonomous vector fields in Hölder classes $C^{\alpha}(\mathbb{T}^{d})$ with $\alpha>\frac{1}{3}$, without any regularity assumptions on the passive scalar except that the initial data is fixed and bounded. The random field has to satisfy very mild conditions in two dimensions, while in higher dimensions it is assumed to have a special geometric structure. The proof relies on dimension-theoretic arguments, as opposed to commutator estimates. A consequence of these results is that anomalous regularization does not occur for this class of random vector fields.
@@ 1. Introduction
The phenomenon of anomalous dissipation, which is the nonvanishing of the viscous energy dissipation in the vanishing viscosity limit, occupies a central position in the mathematical theory of turbulence. It is therefore sometimes referred to as the zeroth law of turbulence. In a seminal 1949 note, Onsager [36] observed that weak solutions of the incompressible Euler equations conserve kinetic energy above Hölder regularity $1/3$, and may fail to conserve kinetic energy below Hölder regularity $1/3$. This prediction, now known as Onsager’s conjecture, has since been resolved: energy conservation holds for $C^{\alpha}$ solutions with $\alpha>1/3$ [30, 17, 28, 15], while non-conservative solutions exist for $\alpha<1/3$ [21, 22, 23, 12, 32, 13]. The exponent $1/3$ thus marks a sharp threshold separating rigidity from flexibility in nonlinear incompressible fluid dynamics. The problem of providing a rigorous example of anomalous dissipation in the vanishing viscosity limit for the (unforced) Navier–Stokes equations in 3D remains open.
A closely related question arises for passive scalar transport. Consider a scalar $\theta:\mathbb{T}^{d}\times[0,T]\rightarrow\mathbb{R}$ which solves the advection–diffusion equation
@@
driven by a divergence-free velocity field $v:\mathbb{T}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ ($\mathbb{T}^{d}$ denotes the $d$-dimensional flat torus; note that in this paper we will consider autonomous vector fields). It is expected by physicists and confirmed by numerical studies that in a turbulent vector field the scalar variance dissipates at a non-vanishing rate as $\varepsilon\to 0$ [37, 26, 38, 39]. Moreover, in the physics literature, beginning with the works of Obukhov, Corrsin, and Yaglom [35, 18, 40] in 1949–1951, dimensional arguments analogous to the Kolmogorov’s 1941 theory of turbulence predict nontrivial flux relations for passive scalars in turbulent regimes. In particular, heuristic cascade theories suggest a relation of the form
$\alpha+2\beta=1,$
linking the Hölder regularity $C^{\alpha}(\mathbb{T}^{d})$ of the velocity field to an effective regularity $C^{\beta}(\mathbb{T}^{d})$ of the scalar. Such predictions are sometimes referred to as manifestations of anomalous regularization, whereby turbulent mixing produces effective smoothing properties that persist in the vanishing diffusivity limit. We will discuss this phenomenon in more detail in Section 4.3.
In mathematics, the existing results only scrape the surface of our understanding of anomalous dissipation or anomalous regularization for passive scalar transport. We are far from providing a rigorous foundation for a phenomenological theory. Only a few deterministic examples of particular vector fields that give rise to anomalous dissipation and occasionally to some version of the Obukhov-Corrsin-Yaglom balance have been built, in [27, 11, 16, 33, 10, 6, 29, 14] (see also [31]), and both anomalous dissipation and anomalous regularization have been recently established in the groundbreaking work [5] for rather general classes of random vector fields with a negative Hölder exponent, close to 1-1. The two effects are partially related as the occurrence of anomalous dissipation implies $\alpha+2\beta\leq 1$, cf. [27, 1]. Some of the aforementioned examples pertain to autonomous vector fields. We do not mention the results in stochastic PDEs pertaining to the Kraichnan model and its relatives as the requirements on roughness in time (white noise) position them too far from the context of the present note.
The present work establishes that in the autonomous setting $\alpha\leq 1/3$ is sharp, in the sense that random Hölder roughness above Onsager’s exponent $1/3$ almost surely prohibits anomalous scalar dissipation or anomalous regularization. We emphasise that we do not make Hölder regularity assumptions on the passive scalar $\theta$, our results can therefore be consider Onsager-supercritical. We consider probability measures on divergence-free autonomous average-free vector fields on the two-dimensional torus satisfying:
  1. (1)
    Hölder regularity $C^{\alpha}(\mathbb{T}^{2})$ with $\alpha>1/3$;
  2. (2)
    a mild nondegeneracy condition ensuring bounded density of the pointwise distribution.
Our main theorem shows that, under these assumptions, the associated transport equation possesses the DiPerna–Lions renormalization property almost surely. The latter property was introduced in the seminal work of DiPerna and Lions [25], where the authors proved it for any vector field with Sobolev regularity $W^{1,1}$. It is a consequence of their theory that then bounded weak solutions of the transport equation are “renormalized” (cf. Definition 2.1), uniqueness holds for the Cauchy problem, and anomalous dissipation is excluded. In fact, for any bounded initial data, solutions of (1.1) converge strongly, as diffusivity vanishes, to the unique inviscid transport solution, and thus scalar dissipation vanishes in the limit. An important extension of the renormalization property to $BV$ vector fields was later achieved by Ambrosio in [3], while the novel Lagrangian approach in [19] has spurred further extensions to other critical regularities, cf. [9]. The sharpness, at the deterministic level, of the thresold $\alpha=1$ in the Sobolev case is known since [24].
Let us state the details.
@@
Consider a bounded divergence-free (autonomous) vector field $v:\mathbb{T}^{d}\rightarrow\mathbb{R}^{d}$. We say that $v$ allows a dissipation anomaly if there is a bounded initial data $\theta_{in}\in L^{\infty}(\mathbb{T}^{d})$ with the following property. If $\theta_{\varepsilon}$ denotes, for any $\varepsilon>0$, the unique bounded solution of the advection-diffusion equation
@@
$\limsup_{\varepsilon\downarrow 0}\,\varepsilon\int_{0}^{T}\int|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt>0$
for some positive $T$.
One main result of this note is the following.
@@
Consider a probability measure $\mathbb{P}$ on the space of continuous divergence-free and average-free autonomous vector fields $v$ on $\mathbb{T}^{2}$. Assume that:
  • (a)
    There is $\alpha>\frac{1}{3}$ such that $v\in C^{\alpha}(\mathbb{T}^{2})$ $\mathbb{P}$-almost surely;
  • (b)
    There is a constant $C>0$ such that
    $\mathbb{P}(\{v:|v(x)-v_{0}|\leq r\})\leq Cr^{2}\qquad\forall r>0,\quad\forall x\in\mathbb{T}^{2}\,,\quad\mbox{and}\quad\forall v_{0}\in\mathbb{R}^{2}\,.$
Then $\mathbb{P}$-almost surely $v$ satisfies the DiPerna-Lions renormalization property (cf. Definition 2.1) and thus does not allow dissipation anomalies.
Note that the condition (b) is extremely mild, cf. Section 4.1, and in fact it can be further relaxed. The assumption that the vector field is average-free almost surely can be also easily removed, cf. Section 4.2: in fact allowing for the addition of a random constant drift only makes the result less interesting, cf. Remark 4.1. Our techniques allow for some extension to higher dimensions, although not quite as clean. We will formulate a general dimensional version in Section 2 below. Here we will limit ourselves to the case of three dimensions because it is easier to state, physically more relevant, and moreover relates the random vector field to the classical “Clebsch variables” of the fluid-dynamics literature.
@@
    There is $\alpha>\frac{1}{3}$ such that $\phi\in C^{1,\alpha}(\mathbb{T}^{3};\mathbb{R}^{2})$ $\mathbb{P}$-almost surely;
@@
Then $\mathbb{P}$-almost surely the vector field $v=\nabla\phi_{1}\times\nabla\phi_{2}$ satisfies the DiPerna-Lions renormalization property and thus does not allow dissipation anomalies.
The DiPerna-Lions renormalization property implies in fact much more, namely:
  • The corresponding transport equation $\partial_{t}\theta+(v\cdot\nabla)\theta=0$ has a unique bounded weak solution $\theta$ for any bounded initial data;
  • $\theta$ is constant on the trajectories of the unique regular Lagrangian flow of $v$ and its $L^{p}$ norms are constant in time;
  • The solutions $\theta_{\varepsilon}$ converge strongly in $C([0,\infty),L^{2}(\mathbb{T}^{d}))$ to $\theta$.
It is from these three properties that one immediately infers
$\lim_{\varepsilon\to 0}\varepsilon\int_{0}^{T}\int|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt=0\,$
for all $T>0$. This consequence of the third property mentioned above is very well-known (in Section 4.4 we give a quick derivation of it from precise statements which can be found in the many lecture notes on the DiPerna-Lions theory available in the literature, e.g. [4, 20]).
The proofs of Theorems 1.2 and 1.3 rely on a probabilistic Morse–Sard mechanism and a striking result of Alberti–Bianchini–Crippa [2] that shows that the renormalization property follows from a geometric condition: if the associated stream function (also referred to as the Hamiltonian) satisfies a Morse–Sard property (its set of critical values has zero Lebesgue measure), then bounded weak solutions are unique and renormalized. The nondegeneracy assumption on the underlying probability law implies, almost surely, a quantitative upper bound on the Hausdorff dimension of the critical set of the stream function. When $\alpha>1/3$, this dimension bound yields a Morse–Sard property via a deterministic covering argument. The theorem of Alberti–Bianchini–Crippa then implies the renormalization property of the vector field.
The exponent $1/3$ arises here from the inequality (cf. Lemma 3.2)
$1+\alpha>2-2\alpha,$
@@
which governs the competition between Hölder regularity and Hausdorff dimension of the critical set. Thus the same threshold that appears in Onsager’s theory for the Euler equations reemerges in a purely geometric mechanism for passive scalar transport. The fact that the threshold exponent is identical is somewhat surprising. First, unlike the nonlinear energy-flux arguments from [17, 30, 28, 15] underlying Onsager’s theorem, the present mechanism is dimension-theoretic. Secondly, arguments supporting the anomalous dissipation for passive scalar transport and the Obukhov-Corrsin-Yaglom law of anomalous regularization, both in physics and in mathematics, rely on the dimensional analysis, energy cascade, and (sometimes) certain statistical homogeneities. To the best of our knowledge, none of them suggests that the exponent $\alpha=1/3$ might be critical in this scenario. In fact, one could argue that the existing deterministic examples, e.g., [27, 16, 33] extending to $\alpha<1$, suggest otherwise. The only exception is [6], which indeed restricts to $\alpha<1/3$, but their example is deterministic as well, and it is hard to say whether it is a hard restriction or an artefact of the method. One should also reiterate that our vector fields are autonomous (but so are the ones in [5] and [33]) and the results are probabilistically a.s., contrary to most of the positive results, with the notable exception of [5], where the authors are close to $\alpha=-1$. The latter work is moreover motivated by preexisting papers in the physics literature which use renormalization group ideas to analyze superdiffusivity, cf. for instance [8] (and also [7, 34]): neither in this literature there seems to be an indication that the exponent $\frac{1}{3}$ (see for instance the Richardson law in [8, p. 216]) is critical. We find it remarkable that the threshold for our arguments in Theorems 1.2 and 1.3 coincides with the Hölder regularity expected for a typical flow in homogeneous turbulence. We do not presently know if this is just a coincidence.
@@ 1.1. Acknowledgments
The research of the first author has been supported by the Cambridge Trust and the Cantab Capital Institute for Mathematics of Information. The research of the second and third authors has been supported by the Simons Foundation through the Simons Initiative on the Geometry of Flows (Grant Award ID BD-Targeted-00017375 CDeL, SM). The third author is also partially supported by the Simons Collaboration on Localisation of Waves, 563916 SM.
2. General theorem and reduction to the Morse-Sard property

### 01 Apr 2026 v2

# Regularity thresholds for anomalous dissipation and related phenomena in passive scalars

@@ Abstract.
We prove the absence of anomalous dissipation for passive scalars driven by some random autonomous divergence-free vector fields in $\mathbb{T}^{d}$. In dimension $d=2$ we just need continuity almost surely and a mild nondegeneracy condition on the randomness. In dimension $d\geq 3$ we assume a special geometric structure and almost sure Hölder regularity with a Hölder exponent bigger than $\frac{1}{8}$. No regularity is assumed on the passive scalar except for boundedness in the initial data. The proof relies on dimension-theoretic arguments, as opposed to commutator estimates. A consequence of these results is that the same assumptions prevent (almost surely) many other expected properties of turbulent flows, such as anomalous regularization, the Yaglom-Obukhov-Corrsin law, and Richardson diffusion.
@@ 1. Introduction
The phenomenon of anomalous dissipation, which is the nonvanishing of the viscous energy dissipation in the vanishing viscosity limit, occupies a central position in the mathematical theory of turbulence. It is therefore sometimes referred to as the zeroth law of turbulence and it is expected to hold for (turbulent) solutions of the classical Navier-Stokes equations as the viscosity converges to zero (or equivalently as the Reynolds number becomes arbitrarily large). A closely related question arises for passive scalar transport. Consider a scalar $\theta_{\varepsilon}:\mathbb{T}^{d}\times[0,T]\rightarrow\mathbb{R}$ which solves the advection–diffusion equation
@@
driven by a divergence-free autonomous velocity field $v:\mathbb{T}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ ($\mathbb{T}^{d}$ denotes the $d$-dimensional flat torus). It is expected by physicists and confirmed by numerical studies that in a turbulent vector field the scalar variance dissipates at a non-vanishing rate as $\varepsilon\to 0$ [51, 30, 52, 53]. This motivates the following definition.
@@
Consider a bounded divergence-free vector field $v:\mathbb{T}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$. We say that $v$ allows a dissipation anomaly if there is a bounded initial data $\theta_{in}\in L^{\infty}(\mathbb{T}^{d})$ with the following property. If $\theta_{\varepsilon}$ denotes, for any $\varepsilon>0$, the unique bounded solution of the advection-diffusion equation
@@
$\limsup_{\varepsilon\downarrow 0}\,\varepsilon\int_{0}^{T}\int|\nabla\theta_{\varepsilon}|^{2}(x,t)\,dx\,dt>0\,.$
Moreover, in the physics literature, beginning with the works of Obukhov, Corrsin, and Yaglom [47, 22, 54] in 1949–1951, dimensional arguments analogous to the Kolmogorov’s 1941 theory of turbulence predict nontrivial flux relations for passive scalars in turbulent regimes. In particular, heuristic cascade theories suggest a relation of the form $\alpha+2\beta=1$, linking the Hölder regularity $C^{\alpha}(\mathbb{T}^{d})$ of the velocity field to an effective regularity $C^{\beta}(\mathbb{T}^{d})$ of the scalar. Such predictions are sometimes referred to as manifestations of anomalous regularization, whereby turbulent mixing produces effective smoothing properties that persist in the vanishing diffusivity limit. We will discuss this phenomenon in more detail in the Sections 3.2 and 3.4. A third related phenomenon is that of Richardson’s dispersion, roughly speaking the assertion that fluid particles get away from each other at a uniform rate independent of their initial position and of the viscosity, see Section 3.3 for the precise definition.
The present work focuses on the case of autonomous vector fields. Under the latter assumption and in two space dimensions we show that all of these phenomena are absent for random continuous vector fields. In $3$ or more space dimensions we are able to draw the same conclusion if we impose some special structure and assume Hölder regularity for the random field above the Hölder exponent $\frac{1}{8}$. The latter looks like a technical point and it seems likely that the special structure assumed in our theorems is an obstruction also at the level of random continuous fields in higher dimensions.
In two space dimensions we consider probability measures on divergence-free autonomous periodic continuous vector fields satisfying a mild nondegeneracy condition, ensured for instance by the absolute continuity of the pointwise distribution. Our main theorem shows that, under these assumptions, the associated transport equation possesses the DiPerna–Lions renormalization property almost surely. The latter property, whose definition dates back to the seminal work of DiPerna and Lions in [29] and is stated precisely in Definition 1.1, seems to obstruct all the expected features of a turbulent flow. This is certainly clear to experts, but since it does not seem to be explicitly stated in a coherent way somewhere, we decided to collect precise statements and give precise proofs in the final part of this paper, without claiming any originality.
Let us now rigorously state the main results.
@@
Consider a probability measure $\mathbb{P}$ on the space of continuous divergence-free autonomous vector fields $v$ on $\mathbb{T}^{2}$. Assume that
$\mathbb{P}(\{v:v(x)=0\})=0\qquad\mbox{for a.e. $x\in\mathbb{T}^{2}$.}$
Then $\mathbb{P}$-almost surely $v$ satisfies the DiPerna-Lions renormalization property (cf. Definition 2.1) and thus does not allow dissipation anomalies (cf. Theorem 3.1), does not support Richardson dispersion (cf. Proposition 3.4), does not display anomalous regularization (cf. Proposition 3.2), and violates Yaglom’s law (under the interpretation of Theorem 3.6).
The main statement, namely that $v$ satisfies the DiPerna-Lions property, follows with a two lines proof: it suffices to apply Fubini’s theorem to reduce it to a powerful deterministic result of Alberti, Bianchini, and Crippa, cf. [3]. In other words, it is rather trivial to see that, even under very mild assumptions, the stream function of a random continuous divergence-free vector field satisfies the “weak Sard property”, cf. Definition 2.2. The latter was introduced in [3] for the more general case of Lipschitz vector field, and the authors showed in fact that it is equivalent to the renormalization property and to the uniqueness of a regular Lagrangian flow in the sense of DiPerna and Lions. In particular it is tempting to advance the following conjecture at the deterministic level:
Conjecture 1.3.
An autonomous average-free and divergence-free vector field $v\in C^{0}(\mathbb{T}^{2},\mathbb{R}^{2})$ allows dissipation anomalies if and only if its stream function violates the weak Sard property of Definition 2.2.
A good reason to believe in the above conjecture is given by the paper [42]. In the latter the authors, inspired by the works relating spontaneous stochasticity and anomalous dissipation (cf. [33]), construct explicit autonomous $2$-dimensional vector fields which allow dissipation anomalies, but the mechanism behind their construction is precisely the nonuniqueness of regular Lagrangian flows.
Following the same blueprint, we can prove some extension of Theorem 1.2 to higher dimensions, although not quite as clean, see Theorem 2.4. Here we will limit ourselves to the case of three dimensions because it is easier to state, physically more relevant, and moreover relates the structure of the random vector field to the classical “Clebsch variables” of the fluid-dynamics literature.
@@
    There is $\alpha>\frac{1}{8}$ such that $\phi\in C^{1,\alpha}(\mathbb{T}^{3};\mathbb{R}^{2})$ $\mathbb{P}$-almost surely;
@@
Then $\mathbb{P}$-almost surely the vector field $v=\nabla\phi_{1}\times\nabla\phi_{2}$ satisfies the DiPerna-Lions renormalization property and hence all the conclusions of Theorem 1.2 apply.
As already mentioned the proofs of Theorems 1.2 and 1.4 rely on a probabilistic Morse–Sard mechanism and a striking result of Alberti–Bianchini–Crippa [3] that shows that the renormalization property follows from a geometric condition. A statement of independent interest which we discovered during our investigations is that the classical Morse-Sard property for (suitable) random maps $\phi:\mathbb{T}^{d}\to\mathbb{R}^{d-1}$ holds, almost surely, with much less regularity than for deterministic maps.
Theorem 1.5.
Consider a probability measure $\mathbb{P}$ on $C^{0}(\mathbb{T}^{d},\mathbb{R}^{d-1})$ and assume that
  • (a)
    $\phi\in C^{1,\alpha}$ almost surely for some $\alpha>\frac{1}{3}$;
  • (b)
    $\mathbb{P}(\{\phi:|D\phi(x_{0})-M_{0}|\leq r\})\leq C_{0}r^{d(d-1)}$ for some fixed $C_{0}$ and all $x_{0}\in\mathbb{T}^{d}$ and $M_{0}\in\mathbb{R}^{d(d-1)}$.
Then $\phi$ has almost surely the classical Morse-Sard property, namely $\phi^{-1}(y)$ is a $1$-dimensional $C^{1}$ submanifold for a.e. $y\in\mathbb{R}^{d-1}$.
1.1. Context of our results
In a seminal 1949 note, Onsager [48] observed that weak solutions of the incompressible Euler equations conserve kinetic energy above Hölder regularity $1/3$, and may fail to conserve kinetic energy below Hölder regularity $1/3$. This prediction, now known as Onsager’s conjecture, has since been resolved: energy conservation holds for $C^{\alpha}$ solutions with $\alpha>1/3$ [37, 21, 34, 19], while non-conservative solutions exist for $\alpha<1/3$ [24, 25, 26, 16, 41, 17]. The exponent $1/3$ thus marks a sharp threshold separating rigidity from flexibility in nonlinear incompressible fluid dynamics. The problem of providing a rigorous example of anomalous dissipation in the vanishing viscosity limit for the (unforced) Navier–Stokes equations in 3D remains open and it seems extremely challenging.
For passive scalars there has been very important progress in recent years, even though the existing results still only scrape the surface of our understanding. In particular the literature is still far from providing a rigorous foundation for a phenomenological theory. Only a few deterministic examples of particular vector fields that give rise to anomalous dissipation and occasionally to some version of the Obukhov-Corrsin-Yaglom balance have been built, in [32, 15, 20, 43, 42, 14, 7, 35, 18] (see also [40]). At the level of random fields, the anomalous dissipation, anomalous regularization, and Richardson dispersion have been recently established for the first time in the groundbreaking work [6] for rather general classes of autonomous random vector fields with a negative Hölder exponent, close to 1-1. It is also proved that anomalous dissipation necessarily implies $\alpha+2\beta\leq 1$, cf. [32, 1]. Some of the aforementioned examples pertain to autonomous vector fields. For the sake of brevity, in this brief overview we will not discuss in detail the results in stochastic PDEs pertaining to the Kraichnan model and its relatives as the requirements on the roughness in time (white noise) on the vector field position them in a different context than the present note, but we refer for instance to [8, 9, 31, 36, 39, 45, 49, 50] (and see references therein).
@@ 1.1. Context of our results
Concerning the thresholds found in this paper, the arguments supporting the anomalous dissipation and the anomalous regularization for passive scalar transport, both in physics and in applied mathematics, rely on the dimensional analysis, energy cascade, and (sometimes) certain statistical homogeneities. To the best of our knowledge, none of them suggests the existence of a critical exponent in the range (1,1)(-1,1), in any dimension. In fact, one could argue that, since most of the existing deterministic examples, e.g., [32, 20, 43] extend to $\alpha<1$, the literature so far suggested otherwise. The only exception is [7], which indeed restricts to $\alpha<1/3$, but their example is also deterministic and it is hard to say whether it is a hard restriction or an artefact of the method. One should also reiterate that our vector fields are autonomous (but so are the ones in [6] and [43]) and the results are probabilistically a.s.. Finally, the work [6] was motivated by preexisting papers in the physics literature which use renormalization group ideas to analyze superdiffusivity and Richardson dispersion, cf. for instance [10] (and also [11, 44]). Neither in this literature there seems to be an indication that some exponent $\alpha<1$ might be critical (see for instance the Richardson law in [10, p. 216]).
@@ 1.2. Acknowledgments
The research of the first author has been supported by the ERC/EPSRC Horizon Europe Guarantee EP/X020886/1 and the by the Istituto Nazionale di Alta Matematica (INdAM), group GNAMPA. The research of the second author has been supported by the Cambridge Trust and the Cantab Capital Institute for Mathematics of Information. The research of the third and fourth authors has been supported by the Simons Foundation through the Simons Initiative on the Geometry of Flows (Grant Award ID BD-Targeted-00017375 CDeL, SM). The fourth author is also partially supported by the Simons Collaboration on Localisation of Waves, 563916 SM.
2. Reduction to the Morse-Sard property
2.1. The DiPerna-Lions property
The following property was introduced in the seminal work of DiPerna and Lions [29], where the authors proved it for any vector field with Sobolev regularity $W^{1,1}$. It is a consequence of their theory that, when all bounded weak solutions of the transport equation are “renormalized”, then uniqueness of the Cauchy problem and strong stability properties hold.
Definition 2.1.
Let $v$ be a bounded divergence-free vector field on $\mathbb{T}^{d}$ (not necessarily autonomous). A bounded weak solution $\theta$ of
$\left\{\begin{array}[]{ll}\partial_{t}\theta+(v\cdot\nabla)\theta=0\\ \\ \theta(\cdot,0)=\theta_{in}\end{array}\right$. (2.1)
is a bounded function $\theta$ which satisfies
$\int_{0}^{\infty}\int_{\mathbb{T}^{d}}\theta(\partial_{t}\varphi+v\cdot\nabla\varphi)\,dx\,dt=\int_{\mathbb{T}^{d}}\theta_{in}(x)\varphi(x,0)\,dx$ (2.2)
The vector field $v$ is said to have the DiPerna-Lions renormalization property if, for every bounded initial data $\theta_{in}$, every bounded weak solution $\theta$ of (2.1), and every test function $\beta\in C^{1}(\mathbb{R})$, the scalar $\beta(\theta)$ is a weak solution of (2.1) with initial data $\beta(\theta_{in})$.
An important extension of the renormalization property to $BV$ vector fields was later achieved by Ambrosio in [4], while the novel Lagrangian approach in [23] has spurred further extensions to other critical regularities, cf. [12]. The sharpness, at the deterministic level, of the threshold $\alpha=1$ in the Sobolev case is known since [28].
2.2. Dimension $d=2$ and the proof of (the renormalization part of) Theorem 1.2
In dimension two the DiPerna-Lions renormalization property for random autonomous divergence-free vector fields follows immediately from a simple Fubini argument and the main result of [3]. In fact we have
$\int\mathcal{L}^{2}(\{v=0\})\,d{\mathbb{P}}(v)=\int_{\mathbb{T}^{2}}{\mathbb{P}}(\{v(x)=0\})\,dx$
and in particular with probability one the zero set of the vector field $v$ has measure zero. Assume now for the moment that the average of $v$ is zero, so that we can introduce the stream function $\phi$ of $v$. We then conclude that the measure
$\mu:=\mathbf{1}_{\{\nabla\phi=0\}}\mathcal{L}^{2}$
vanishes identically and in particular that $\phi$ satisfies the weak Sard property in [3], which is the condition $\phi_{\sharp}\mu\perp\mathcal{L}^{1}$ (see Definition 2.2). This is then proved in [3] to be equivalent to the DiPerna-Lions renormalization property. As already mentioned, the derivation of the further conclusions in Theorem 1.2 will be given in Sections 3.1-3.4.
If $v$ does not have average 0, then there is no global stream function for $v$ on the torus. However a stream function exists in every subset $\Omega$ which is simply connected. The result in [3] can then be used to show that, when $\Omega$ is simply connected and $\theta$ and $\beta$ are as in Definition 2.1, then $\beta(\theta)$ will be a distributional solution of the transport equation (2.1) in $\Omega\times[0,\infty)$ with initial data $\beta(\theta_{in})$. But a simple partition of unity allows then to show that $\beta(\theta)$ is in fact a distributional solution on the whole torus.
Observe that the argument for the weak Sard property are not really dimension dependent, namely the following holds.
Definition 2.2.
Let $\phi\in C^{1}(\mathbb{T}^{d},\mathbb{R}^{d-1})$ and let $Z$ be its critical set $Z=\{x:{\rm rank}\,(D\phi(x))\leq d-2\}$. Then we say that $\phi$ has the weak Sard property if $\phi_{\sharp}(\mathbf{1}_{Z}\mathcal{L}^{d})\perp\mathcal{L}^{d-1}$.
Proposition 2.3.
Consider a probability measure on the space $C^{1}(\mathbb{T}^{d},\mathbb{R}^{d-1})$ such that $\mathbb{P}(\{\phi:{\rm rank}\,D\phi(x_{0})<d-1\})=0$ for every $x_{0}$. Then its critical set $Z$ is, $\mathbb{P}$-almost surely, a Lebesgue null set and therefore $\phi$ has the weak Sard property.
2.3. Higher dimensions